OronSH:

A [i]Nim-style[/i] game is defined as follows. Two positive integers \(k\) and \(n\) are specified, along with a finite set \(S\) of \(k\)-tuples of integers (not necessarily positive). At the start of the game, the \(k\)-tuple \((n, 0, 0, ..., 0)\) is written on the blackboard. A legal move consists of erasing the tuple \((a_1,a_2,...,a_k)\) which is written on the blackboard and replacing it with \((a_1+b_1, a_2+b_2, ..., a_k+b_k)\), where \((b_1, b_2, ..., b_k)\) is an element of the set \(S\). Two players take turns making legal moves, and the first to write a negative integer loses. In the event that neither player is ever forced to write a negative integer, the game is a draw. Prove that there is a choice of \(k\) and \(S\) with the following property: the first player has a winning strategy if \(n\) is a power of 2, and otherwise the second player has a winning strategy.

minesweeper:

this is trivial

minesweeper:

trivial because alecks bad

OronSH:

couldn't make mop you can only mop the floors

minesweeper:

so bad imagien making mops

minesweeper:

you could only make mops

minesweeper:

bad!!!!!!38>37

minesweeper:

ok but when 7<8

minesweeper:

every time

minesweeper:

bruh but imagien \(11^{\frac{1}{7}}>13^{\frac{1}{8}}\)

OronSH:

where is your solution

OronSH:

please solve the 2015 tstst problem 6 or bad

minesweeper:

We have $k=9$, and $S$ consists of $S_1=$ (-1,0,0,1,0,0,0,0,0) and (-1,1,0,-1,0,0,0,0,0) $S_2=$ (0,-1/-2/-3,0,-1,0,0,1,0,0) and (0,-1,1,0,-1,0,-1,0,0) $S_3=$ (-1,0/-1/-2,0,0,-1,0,0,0,0) and (-2,1,0,0,-1,0,0,0,0) $S_4=$ (0/-1,0,0,-1,-1,0,0,0,0) $S_5=$ (-1,0/-1/-2,0,0,0,0,-1,0,0) and (-2/-3,0,0,0,0,0,-1,0,0) $S_6=$ (-1/-2,0,0,-2,0,0,0,0,0) and (0,0/-1/-2,0,-2,0,0,0,0,0) and for each of the above 9-tuple also take the two other "cyclic within triplet" permutations, i.e.: (123456789) $\rightarrow$ (312645978) and (231564897). The / above means "and". So for example $S_4$ consists of 6 elements (0,0,0,-1,-1,0,0,0,0); (-1,0,0,-1,-1,0,0,0,0); (0,0,0,0,-1,-1,0,0,0); (0,-1,0,0,-1,-1,0,0,0); (0,0,0,-1,0,-1,0,0,0); (0,0,-1,-1,0,-1,0,0,0). more intuitive description of each set Firstly we note that (1,0,0...) is winning, since P1 (player 1) can play the first k-tuple in $S_1$ to reach (0,0,0,1,0...), and P2 has no legal move. Next we proceed inductively in a few parts. In all inductions we induct by decreasing the total sum of all elements in the k-tuple. Note that since our construction is "cyclic within triplet", all induction statements are also true for "cyclic within triplet" permutations. Part 1: (a,b,0,0...) is losing for player 1 when a is odd, except for when (a,b) = (1,0). Proof: All the k-tuples in $S$ decrease some of the 4th to 9th number in the k-tuple except for those in $S_1$. So P1 only has two possible choices: Case 1: He goes to (a-1,b,0,1,0...). If $a\ge 3$ P2 then plays from $S_1$ to reach (a-2,b+1,0,0...) and we are done by induction. If $a=1$ then P2 plays from the first k-tuple in $S_2$ to reach (0,c,0,0,0,0,1,0,0) where c<b is either 0 or any odd number $\ge 3$ (possible as P2 can choose to decrease b by 1,2 or 3, and $b\neq 0$ as a=1). Now P1 has lost when c=0 & otherwise can only play from $S_1$ to reach (0,c-1,0,0,1,0,1,0,0). P2 plays from 2nd k-tuple in $S_2$ to reach (0,c-2,1,0,0...) and by induction this is losing. Case 2: He goes to (a,b-1,0,0,1,0...). If $a=1$ P2 plays from the first k-tuple in $S_3$ to reach (0,c,0,0...) where c<b is either 0 or any odd number $\ge 3$, and by induction this losing. If $a\ge 3$ then P2 plays from the second k-tuple in $S_3$ to reach (a-2,b,0,0...) and we are done by induction (note b$\neq 0$ is implied in case 2). Part 2: (a,b,0,0...) is losing for player 1 when a is even, $a\ge 2$ and $a+2b$ is not a power of 2. Proof: Similarly, we know that P1 must play some k-tuple in $S_1$, so there are two possible options: Case 1: He goes to (a-1,b,0,1,0...). P2 plays from $S_1$ again to reach (a-2,b+1,0,0...). If $a\ge 4$ we are done by induction as a+2b remains not a power of 2. If $a=2$ and b+1 is odd we are done by part 1 (note $b\neq0$ as a+2b not power of 2), while when b+1 is even we know that $\frac{a+2b}2=b+1$ is also not a power of 2 and we are done by induction. Case 2: He goes to (a,b-1,0,0,1,0...). P2 plays from first k-tuple in $S_3$ to reach (a-1,b-1,0,0...) and by part 1 we are done (note (a,b) is not (2,1) as that gives a+2b=4). Part 3: (a,b,0,0...) is winning for player 1 when a is even, $a\ge 2$ and $a+2b$ is a power of 2. Proof: P1 plays from $S_1$ to reach (a-1,b,0,1,0...). P2 has four options. Case 1: He plays from $S_1$ to reach (a-2,b+1,0,0...). We are done by induction since either $a\ge 4$ and a+2b remains a power of 2, or a=2 and b+1 is a power of 2 (b=0 reduces to trivial base case as well) Case 2: He plays from $S_1$ to reach (a-1,b-1,0,1,1,0...). P1 now plays from $S_4$ to reach (a-1,b-1,0,0...) and by part 1 we are done. If a=2,b=1, player 1 goes to (a-2,b-1,0,0...) instead. Case 3: He plays from $S_1$ to reach (a-2,b,0,2,0...). If $a\ge 4$ P1 plays from first k-tuple of $S_6$ to reach (a-3,b,0,0...) and we are done by part 1, except when a=4,b=0 where he goes to (a-4,b,0,0...) instead. If $a=2$ he plays from second k-tuple in $S_6$ to reach (0,c,0,0...) where c$\le$b is either 0 or an odd number $\ge 3$ and we are done by part 1. Case 4: He plays from $S_2$ to reach (a-1,b-1/2/3,0,0,0,0,1,0,0). If $a\ge 4$ P1 plays from the second k-tuple in $S_5$ to reach (a-3,b-1/2/3, 0,0...) and we are done by part 1, except when a=4 and b-1/2/3=0, where he goes to (0,0,0...) instead. If $a=2$ P1 plays from the first k-tuple in $S_5$ to reach (0,c,0,0...) where c<b is either 0 or an odd number $\ge 3$ and we are done by part 1.

OronSH:

sobad not even \(\LaTeX\)

Fortish:

@oronsh wrote:
sobad not even \(\LaTeX\)
May not have \(\LaTeX\) but he tried to help, thats the point. It doesn't have to be perfect anyways.

OronSH:

also he copy pasted the solution from aops lol imagine not even solving 55 mohs so bad

minesweeper:

alecks bad

OronSH:

oren bad you couldn't even copy paste the latex correctly